# Hilbert property for low-genus families and degree-one del Pezzo surfaces

### v1

@@ 3. Low-genus families on degree-one del Pezzo surfaces
Let $f(t)\in K[t]$ be a cubic polynomial and define
@@
    and the induced map $S^{\text{aff}}\to W^{\text{aff}}$ is also flat by preservation of flatness under base-change. Now, any variety $V$ with a flat surjective morphism $V^{\prime}\to V$ from a smooth variety $V^{\prime}$ is smooth (see [18, Thm. 23.7]), hence $W^{\text{aff}}$ is smooth. Then all singularities of $W$ lie on $X_{3}=0$.
@@
    To determine the singularity type/types, we follow [2, §2].
@@
        Since $G_{W}(0,0,1)\neq 0$, one of the singularities is $A_{2}$ by [2, Lem. 3], hence so is the other singularity by symmetry.
@@
      Since $g_{1}(X_{1}):=G_{W}(0,X_{1},1)$ has a zero of order $1$ at $X_{1}=0$ and $g_{0}(X_{0}):=G_{W}(X_{0},0,1)$ has a zero of order $3$ at $X_{0}=0$, the singularity is $A_{5}$ [2, Lem. 3].
@@
For each point $P=[x_{0}:y_{0}:z_{0}:w_{0}]\in S$ with $w_{0}\neq 0$, the set
$\{R\in S:|C_{R}\cap D_{P}|<3\}$
is a proper closed subset of $S$.
@@
The curve $D_{P}$ is irreducible, thus so is $B_{P}$, which coincides with the plane section of $W\subset\mathbb{P}^{3}$ given by $\Pi_{P}:w_{0}^{3}X_{2}-w_{0}^{3}f(z_{0}/w_{0})X_{3}=0$. Let $L\subset\Pi_{P}$ be a general line. This intersects $B_{P}$ transversely in 3 points. Let $\sigma:W^{\prime}\rightarrow W$ be the composition of the minimal desingularization of $W$ and the blowup in the inverse image of $L\cap W$. Let $g:W^{\prime}\rightarrow\mathbb{P}^{1}$ be the genus-$1$ fibration corresponding to the pencil of plane sections cut out in $W$ by planes containing $L$.
Any relatively minimal model of $W^{\prime}$ is a rational elliptic surface, which has (étale/topological) Euler number $12$ [23, Ch. 5], making the Euler number $e$ of $W^{\prime}$ at least $12$. The Euler number of a regular genus-$1$ fibration is the sum $\sum_{v}e(F_{v})$ of the Euler numbers of its singular fibers $F_{v}$ (loc. cit.). Consider now the planes $\Pi_{\pm}$ passing through $L$ and $0:\pm\sqrt{c}:1:0]$, with the convention that $\Pi_{+}=\Pi_{-}$ in the case $c=0$. Let $F_{P},F_{\pm}$ be the fibers corresponding to the plane sections $\Pi_{P}\cap W$ and $\Pi_{\pm}\cap W$, respectively.
@@
We know that $B_{P}$ is irreducible. We claim that each $\Pi_{\pm}\cap W$ is also irreducible. In fact, since $L$ is a general line on $\Pi_{P}$, it is skew with respect to any finite set of lines in $\mathbb{P}^{3}$ not lying on $\Pi_{P}$. There are only finitely many lines on $W$ since it is not a cone (Lemma 3.3), and none of these lies on $\Pi_{P}$ since $\Pi_{P}\cap W=B_{P}$ is irreducible, hence $L$ is skew with respect to all lines of $W$. Then $\Pi_{\pm}\cap W$ are plane cubics with no lines, proving the claim.
From irreducibility above, we infer that $F_{P}\cong\Pi_{P}\cap W$ has $1$ component, while each $F_{\pm}$ has at most $3$ components when $F_{+}\neq F_{-}$ (equiv. $c\neq 0$) and $6$ or $7$ components when $F_{+}=F_{-}$ (equiv. $c=0$): the proper transform of $\Pi_{\pm}\cap W$ and the exceptional components coming from the desingularization (the numbers come from Lemma 3.3: $A_{k}$ and $E_{k}$-singularities give $k$ exceptional components). In each case we have $\sum_{F_{v}\in\{F_{+},F_{-}\}}e(F_{v})\leq 8$, thus $\sum_{F_{v}\in\{F_{P},F_{+},F_{-}\}}e(F_{v})\leq 10$, and there is at least one other singular fiber $F_{Q}\neq F_{P},F_{\pm}$.
Now $\sigma(F_{Q})$ is a plane section of $W$ with some singular point $R^{\prime}\neq[0:\pm\sqrt{c}:1:0]$, hence $\sigma(F_{Q})=T_{R^{\prime}}W\cap W$. Then, for $R\in\theta^{-1}(R^{\prime})$, the curve $C_{R}=\theta^{-1}(\sigma(F))\subset S$ intersects $D_{P}$ in at least 3 points. Since $C_{R}\cdot D_{P}=3$, we are done. ∎
@@
Denote by $\pi_{1}:\mathcal{C}\rightarrow S^{o},\pi_{2}:\mathcal{C}\rightarrow S$ the two projections. Define the normalization morphism $r:\mathcal{C}^{\text{n}}\to\mathcal{C}$, and let $\pi^{\text{n}}_{i}:=\pi_{i}\circ r$ for $i=1,2$ be the composition.
@@
The projection $\pi_{1}:\mathcal{C}\rightarrow S^{o}$ is flat and $\mathcal{C}$ is geometrically integral.
@@
The fibers of $\pi_{1}$ are curves in $|{-3K_{S}}|$, and in particular they all have the same Hilbert polynomial (with respect to any fixed ample line bundle on $S$), making $\pi_{1}$ flat [14, Thm. III.9.9]. For general $P$, the curve $C_{P}$ is geometrically integral [20, Prop. 3.16]. The morphism $\pi_{1}:\mathcal{C}\to S^{o}$ is now flat over a geometrically integral base with geometrically integral generic fiber, hence $\mathcal{C}$ is geometrically integral [17, Prop. 4.3.8]. ∎
Recall the blow-up map $\rho\colon\mathcal{E}\rightarrow S$.
@@
The proof of [20, Thm. 3.1] shows that, for infinitely many $R\in\mathbb{P}^{1}(K)$, the surface $\mathcal{C}_{R}:=\mathcal{C}\times_{\pi_{1},S}D_{R}$ (i.e. the restriction of $\mathcal{C}$ to $P\in D_{R}$) has dense $K$-rational points. In particular, $\{R\in\mathbb{P}^{1}(K):(\pi\circ\rho^{-1}\circ\pi_{1})^{-1}(R)(K)\text{ is dense in }(\pi\circ\rho^{-1}\circ\pi_{1})^{-1}(R)\}$ is dense in $\mathbb{P}^{1}(K)$, thus $\mathcal{C}(K)$ is dense in $\mathcal{C}$. ∎
@@
The morphism $\pi_{2}:\mathcal{C}\rightarrow S$ has geometrically integral generic fiber.
@@
Assume that $\pi_{2}$ does not have a geometrically integral generic fiber; then the same holds for $\pi^{\text{n}}_{2}$, and we get a relative normalization decomposition:
@@
First extension. The map $\pi_{1}$ admits a rational section $S\dashrightarrow\mathcal{C},\ P\mapsto(P,Q)$ with $Q\in D_{P}$ such that Q=[2]PQ=[-2]P on the elliptic curve $D_{P}$ with origin $\mathcal{O}$ (see Remark 3.5). For general $P$, this point $Q$ has multiplicity $1$ in the intersection $D_{P}\cap C_{P}$ and is thus smooth in $C_{P}$, hence $(P,Q)$ lies in the smooth locus of $\mathcal{C}$. In particular, this section lifts to a map $s:S\dashrightarrow\mathcal{C}^{\text{n}}$. The composition $\pi^{\text{n}}_{2}\circ s:S\dashrightarrow S$ is the map $2]:S\dashrightarrow S$, and thus $K(S^{\prime})$ is contained in the degree-$4$ extension $K(S)_{2}/K(S)$ (with $K(S)_{2}\cong K(S)$ as abstract fields) given by multiplication by [2][-2].
Second extension. For general $P$, the restriction $\mathcal{C}^{n}_{P}:=\mathcal{C}^{n}\times_{\mathcal{C}}C_{P}\to C_{P}$ of the normalization morphism $\mathcal{C}^{n}\to\mathcal{C}$ above $C_{P}\cong\{P\}\times C_{P}\subset\mathcal{C}$ is the normalization morphism of $C_{P}$ and $C_{P}$ has a simple node at $P$, thus the inverse image $V$ in $\mathcal{C}^{n}$ of a sufficiently small Zariski-open $U$ of the diagonal $\Delta:=\{(P,P):P\in S^{o}\}\subset\mathcal{C}$ is an étale double cover $V\to U$. The composition
$V\to\mathcal{C}^{\text{n}}\xrightarrow{\pi^{\text{n}}_{2}}S$
is a generically finite map of degree $2$, and by (1) the field extension $K(S^{\prime})/K(S)$ of degree $\geq 2$ embeds in the étale extension $K(V)/K(S)$ of degree $2$. Hence $K(S^{\prime})=K(V)$.
Linear disjointness. Since the Mordell–Weil group of $\mathcal{E}$ is torsion-free [23, Thm. 7.4], the degree-$4$ extension $K(S)_{2}/K(S)$ given by multiplication by [2][-2] has no non-trivial proper subextensions, as any of these would have degree $2$ and would give rise to a cyclic 2-isogeny on the elliptic surface $\mathcal{E}$, hence to a non-trivial $2$-torsion section. This immediately implies linear disjointness and concludes the proof. ∎
We now prove that the surface $S$ satisfies the Hilbert property.
@@
Consider the low-genus family $\pi^{\text{n}}_{2}:\mathcal{C}^{\text{n}}\to S$ with underlying fibration $\pi^{\text{n}}_{1}\colon\mathcal{C}^{\text{n}}\rightarrow S^{o}$ (this is a family with general fiber the normalization of some curve $C_{P}$ and general element $C_{P}$). We let $X$ be its inverse image in $\mathcal{E}$, i.e. $X=\mathcal{C}^{\text{n}}\times_{\pi_{2}^{\text{n}},\rho}\mathcal{E}$. The variety $X$ has a low-genus fibration with base $S^{o}$, namely the composition $X\to\mathcal{C}^{\text{n}}\xrightarrow{\pi_{1}^{\text{n}}}S^{o}$. We denote by $p_{2}$ the second projection $X\to\mathcal{E}$. By Proposition 3.10, $\pi^{\text{n}}_{2}$ has geometrically integral generic fiber, and thus the surjective base change $p_{2}$ of $\pi_{2}^{\text{n}}$ has this also.
Let $F$ denote the union of divisors on $\mathcal{E}$ which are vertical with respect to the elliptic fibration $\pi\colon\mathcal{E}\rightarrow\mathbb{P}^{1}$ and do not intersect $\rho^{-1}(C_{P})\subset\mathcal{E}$ for a general $P$ transversally. As the fibers of $\pi$ are all geometrically integral (they are Weierstrass curves), the irreducible divisors in $F$ are fibers of $\pi$. By Proposition 3.6, the only such fiber that possibly intersects the curve $C_{P}$ non-transversally for a general $P$ is the fiber $\pi^{-1}([1:0])$ at infinity. So $\mathcal{E}\setminus F$ contains $\pi^{-1}(\mathbb{A}^{1})\cong S\setminus Z$ for $Z$ the anticanonical divisor of $S$ given by $w=0$. This is simply connected, as follows for instance by applying a result of Coccia [3, Prop. 4.3(i)] (see also [14, Lem. 3.27] for the case $a_{6}(0)\neq 0$, or equivalently, smooth ZZ). [^1] Following a similar approach to Coccia’s, one may also prove this simple connectedness by first noting that the fundamental group of $S\setminus Z\cong\pi^{-1}(\mathbb{A}^{1})$ is abelian using e.g. [13, X.1.4]. Over an algebraically closed field of characteristic 0, a smooth variety with abelian fundamental group is simply connected if and only if its Picard group is torsion-free and it has no non-constant invertible functions. As the anticanonical class $[Z]$ is primitive in the torsion-free module $\operatorname{Pic}S$, both conditions hold for $S\setminus Z$.
The existence of a $P$ as in the statement gives Zariski–density of $S(K)$ by Nijgh’s work [20], and the Hilbert property for $S$ follows from Theorem 1.1. ∎
@@
Let $U\subset\mathbb{A}_{K}^{9}$ be the nonempty Zariski-open whose $\bar{K}$-points are $9$-tuples $(a,b,\ldots,f_{3})$ corresponding to a smooth $S$, and let $X\subset U\times\mathbb{P}(2,3,1,1)$ be the total space associated to the parametric family of Definition 3.1. Then the projection $\pi=pr_{1}:X\to U$ is smooth surjective and its fibers are del Pezzo surfaces of degree $1$. Note that $X$ is birational to the affine hypersurface $Y$ in $\mathbb{A}_{K}^{11}$ defined by the equation
$y^{2}=x^{3}+(af(z)+b)x+(cf(z)^{2}+df(z)+e).$
The hypersurface $Y$ is rational via projection to any coordinate appearing linearly.
@@
For any $x\in V_{N}(K)$, the del Pezzo surface $\pi^{-1}(\pi(x))$ has the Hilbert property by Theorem 1.2. The morphism $\pi|_{V_{N}}:V_{N}\to U$ has geometrically integral generic fiber and thus sends non-thin sets to non-thin sets. Then, since $V_{N}(K)$ is non-thin by Hilbert’s Irreducibility Theorem (recall that $X$, and thus $V_{N}$, is rational), $\pi(V_{N}(K))$ is non-thin.

### v2

@@ 3. Low-genus families on degree-one del Pezzo surfaces
Let $f(t)=f_{3}t^{3}+f_{2}t^{2}+f_{1}t+f_{0}\in K[t]$ be a cubic polynomial and define
@@
    and the induced map $S^{\text{aff}}\to W^{\text{aff}}$ is also flat by preservation of flatness under base-change. Now, any variety $V$ with a flat surjective morphism $V^{\prime}\to V$ from a smooth variety $V^{\prime}$ is smooth (see [23, Thm. 23.7]), hence $W^{\text{aff}}$ is smooth. Then all singularities of $W$ lie on $X_{3}=0$.
@@
    To determine the singularity type/types, we follow [3, §2].
@@
        Since $G_{W}(0,0,1)\neq 0$, one of the singularities is $A_{2}$ by [3, Lem. 3], hence so is the other singularity by symmetry.
@@
      Since $g_{1}(X_{1}):=G_{W}(0,X_{1},1)$ has a zero of order $1$ at $X_{1}=0$ and $g_{0}(X_{0}):=G_{W}(X_{0},0,1)$ has a zero of order $3$ at $X_{0}=0$, the singularity is $A_{5}$ [3, Lem. 3].
@@
We give an explicit description of the ramification locus of $\theta$:
Lemma 3.4.
The ramification locus of $\theta$ is the union of the two (possibly coinciding) points $0:\pm f_{3}\sqrt{c}:1:0]$ and the two (possibly coinciding) divisors $\{z=\rho_{1}w\},\{z=\rho_{2}w\}$, where $\rho_{1}$ and $\rho_{2}$ denote the roots of $f^{\prime}(t)$.
Here $f^{\prime}(t)=3f_{3}t^{2}+2f_{2}t+f_{1}$ indicates the derivative of $f$. The points $0:\pm f_{3}\sqrt{c}:1:0]$ are the points of $S$ lying above the singular points $0:\pm\sqrt{c}:1:0]$ of $W$.
Proof.
The restriction of $\theta$ to $S^{\text{aff}}$ is the restriction of the map $\phi:\mathbb{A}^{3}\to\mathbb{A}^{3}:(x,y,z)\mapsto(x,y,f(z))$. The map $\phi$ ramifies precisely when $f^{\prime}(z)=0$, i.e. on the two planes defined by $z=\rho_{1},z=\rho_{2}$. Since the ramification locus of a morphism is invariant under arbitrary base-change (the ramification locus is the support of the sheaf of relative differentials, so this follows from the fact that the latter is invariant under arbitrary base-change [2, Tag 01V0]), and the morphism $S^{\text{aff}}\to W^{\text{aff}}$ is the base-change of $\phi$ along the inclusion $W^{\text{aff}}\subset\mathbb{A}^{3}$, the ramification locus of $\theta|_{S^{\text{aff}}}=\phi|_{S^{\text{aff}}}:S^{\text{aff}}\to W^{\text{aff}}$ are then the two plane sections defined by $z=\rho_{1},z=\rho_{2}$ of $S^{\text{aff}}$. Thus the ramification locus of $\theta$ contains their projective closures $\{z=\rho_{1}w\},\{z=\rho_{2}w\}$.
Let now $D:\{w=0\}$ be the divisor at infinity of $S$. It remains to determine on which points of $D$ the map $\theta$ ramifies. Looking at the expression for $\theta$, we see that $D$ is the scheme-theoretic inverse image $\theta^{-1}(l)$ of the line $l:\{X_{0}=X_{3}=0\}$. Hence the morphism
$\alpha=\theta|_{D}:D\to l,\ [x:y:z:0]\mapsto[0:y:f_{3}z^{3}:0]$
is the base-change of $\theta$ along the inclusion $l\subset W$. Again applying preservation of the ramification locus under base-change, the points of $D$ where $\theta$ ramifies coincide with those where $\alpha$ ramifies. The point $[1:1:0:0]$ is contained in both ramification divisors listed above, and so it is a ramified point. The complement $D\setminus\{[1:1:0:0]\}$ is fully contained in the chart $z\neq 0$, where $D$ restricts to the affine Weierstrass curve $y^{2}=x^{3}+cf_{3},w=0$, and $\alpha$ restricts to the projection $(x,y,0)\mapsto(0,y,0)$, which ramifies precisely at the points $(0,\pm f_{3}\sqrt{c},0)$, giving the points $0:\pm f_{3}\sqrt{c}:1:0]$ in the weighted projective space. ∎
@@
In view of our wish to apply Theorem 1.1 to the fibration $\pi$ and the family $\{C_{P}\}_{P\in S^{o}}$, we show that the general member of the family intersects the curves of the pencil $|{-K_{S}}|$ transversally.
@@
For each $D\in|{-K_{S}}|$, the set
$\{R\in S^{o}:|C_{R}\cap D|<3\}$
is a proper closed subset of $S^{o}$.
@@
We treat first the case where $D$ is not the divisor $\{w=0\}\in|{-K_{S}}|$. The curve $D$ is irreducible, thus so is $B=\theta(D)$. This image $B$ coincides with the plane section of $W\subset\mathbb{P}^{3}$ given by $\Pi:w_{0}^{3}X_{2}-w_{0}^{3}f(z_{0}/w_{0})X_{3}=0$.
Let $L\subset\Pi$ be a general line. This intersects $B$ transversely in 3 points. Let $\sigma:W^{\prime}\rightarrow W$ be the composition of the minimal desingularization of $W$ and the blowup in the inverse image of $L\cap W$. Let $g:W^{\prime}\rightarrow\mathbb{P}^{1}$ be the genus-$1$ fibration corresponding to the pencil of plane sections cut out in $W$ by planes containing $L$.
Any relatively minimal model of $W^{\prime}$ is a rational elliptic surface, which has (étale/topological) Euler number $12$ [27, Ch. 5], making the Euler number $e$ of $W^{\prime}$ at least $12$. The Euler number of a regular genus-$1$ fibration is the sum $\sum_{v}e(F_{v})$ of the Euler numbers of its singular fibers $F_{v}$ (loc. cit.). Consider now the planes $\Pi_{\pm}$ passing through $L$ and $0:\pm\sqrt{c}:1:0]$, with the convention that $\Pi_{+}=\Pi_{-}$ in the case $c=0$. Let $F,F_{\pm}$ be the fibers corresponding to the plane sections $\Pi\cap W$ and $\Pi_{\pm}\cap W$, respectively.
@@
We know that $B$ is irreducible. We claim that each $\Pi_{\pm}\cap W$ is also irreducible. In fact, since $L$ is a general line on $\Pi$, it is skew with respect to any finite set of lines in $\mathbb{P}^{3}$ not lying on $\Pi$. There are only finitely many lines on $W$ since it is not a cone (Lemma 3.3), and none of these lies on $\Pi$ since $\Pi\cap W=B$ is irreducible, hence $L$ is skew with respect to all lines of $W$. Then $\Pi_{\pm}\cap W$ are plane cubics with no lines, proving the claim.
From irreducibility above, we infer that $F\cong\Pi\cap W$ has $1$ component, while each $F_{\pm}$ has at most $3$ components when $F_{+}\neq F_{-}$ (equiv. $c\neq 0$) and $6$ or $7$ components when $F_{+}=F_{-}$ (equiv. $c=0$): the proper transform of $\Pi_{\pm}\cap W$ and the exceptional components coming from the desingularization (the numbers come from Lemma 3.3: $A_{k}$ and $E_{k}$-singularities give $k$ exceptional components). In each case we have $\sum_{F_{v}\in\{F_{+},F_{-}\}}e(F_{v})\leq 8$, thus $\sum_{F_{v}\in\{F,F_{+},F_{-}\}}e(F_{v})\leq 10$, and there is at least one other singular fiber $F_{Q}\neq F,F_{\pm}$.
Now $\sigma(F_{Q})$ is a plane section of $W$ with some singular point $R^{\prime}\neq[0:\pm\sqrt{c}:1:0]$, hence $\sigma(F_{Q})=T_{R^{\prime}}W\cap W$. Then, for $R\in\theta^{-1}(R^{\prime})$, the curve $C_{R}=\theta^{-1}(\sigma(F))\subset S$ intersects $D$ in at least 3 distinct points. Since $C_{R}\cdot D=3$, we are done.
Finally, let us consider the case $D=\{w=0\}$. In this case, looking at the expression for $\theta$ we see that $\theta(D)$ is the line $l=\{X_{0}=X_{3}=0\}$ and that $D=\theta^{-1}(l)$ (as already noticed in the proof of Lemma 3.4). Take now any $R$ such that $\theta(R)$ does not lie on any of the finitely many lines on $W$ passing through its singular points $\{[0:\pm\sqrt{c}:1:0]\}$. With this choice of $R$, the plane $T_{\theta(R)}W$ does not contain any of the points $\{[0:\pm\sqrt{c}:1:0]\}$. Indeed, if it did, then the line passing through $\theta(R)$ and this point would intersect $W$ with multiplicity $\geq 4$, with a contribution of at least $2$ both from $\theta(R)$ and the singular point.
We conclude by observing that the inverse image $C_{R}=\theta^{-1}(T_{\theta(R)}W\cap W)=\theta^{-1}(T_{\theta(R)}W)$ intersects $D=\theta^{-1}(l)$ in the subscheme $\theta^{-1}(l\cap T_{\theta(R)}W)$, and the latter has at least $3$ distinct geometric points by Lemma 3.4. ∎
Although not needed in this paper, we include the following for completeness:
Lemma 3.8.
For general $P$, the curve $C_{P}$ has geometric genus $1$.
Proof.
Let $P\in S(\bar{K})$ be general. We first observe that, as argued at the end of Proposition 3.7, the tangent plane $T_{\theta(P)}W$ does not contain any of the singular points of $W$. Thus, apart from the three simple nodal singularities in $\theta^{-1}(P)$, the only other singularities the inverse image $\theta^{-1}(T_{\theta(P)}W\cap W)$ can acquire are at the points of non-transverse intersection of the curve $T_{\theta(P)}W\cap W$ with the branch locus of $\theta$ (at transverse intersection points the inverse image is smooth, e.g. this can be seen using the aforementioned argument presented in [8, p. 595]).
Lemma 3.4 tells us that the branch locus $B$ of $\theta$ is the union $\theta(\{z=\rho_{1}w\})\cup\theta(\{z=\rho_{2}w\})\cup W^{\text{sing}}$, where $\rho_{1}$ and $\rho_{2}$ denote the roots of $f^{\prime}(t)=0$. The divisors $\{z=\rho_{i}w\}_{i=1,2}$ are part of the pencil $|{-K_{S}}|$ and neither of them is the curve $\{w=0\}\in|{-K_{S}}|$. The argument of the first part of the proof of Proposition 3.7 gives that the curve $T_{\theta(P)}W\cap W$ is transverse to both of them, concluding the proof. ∎
We introduce the total space $\mathcal{C}$ for the family $\{C_{P}\}_{P\in S^{o}}$.
@@
Denote by $p:\mathcal{C}\rightarrow S^{o}$ and $\phi:\mathcal{C}\rightarrow S$ the two projections. Define the normalization morphism $\nu:\mathcal{C}^{\text{n}}\to\mathcal{C}$, and let $p^{\text{n}}:=p\circ\nu,\phi^{\text{n}}:=\phi\circ\nu$ be the two compositions.
@@
The projection $p:\mathcal{C}\rightarrow S^{o}$ is flat and $\mathcal{C}$ is geometrically integral.
@@
The fibers of $p$ are the members of the family $\{C_{P}\}_{P\in S^{o}}$. These curves lie in $|{-3K_{S}}|$, and in particular they all have the same Hilbert polynomial (with respect to any fixed ample line bundle on $S$), making $p$ flat [18, Thm. III.9.9]. For general $P$, the curve $C_{P}$ is geometrically integral [24, Prop. 3.16]. The morphism $p:\mathcal{C}\to S^{o}$ is now flat over a geometrically integral base with geometrically integral generic fiber, hence $\mathcal{C}$ is geometrically integral [21, Prop. 4.3.8]. ∎
Recall the blow-up map $\rho\colon\mathcal{E}\rightarrow S$. As already sketched in the introduction, the works [13] and [24] give:
@@
The proof of [24, Thm. 3.1] shows that, for infinitely many $R\in\mathbb{P}^{1}(K)$, the surface $\mathcal{C}_{R}:=\mathcal{C}\times_{p,S}D_{R}$ (i.e. the restriction of $\mathcal{C}$ to $P\in D_{R}$) has dense $K$-rational points. In particular, $\{R\in\mathbb{P}^{1}(K):(\pi\circ\rho^{-1}\circ p)^{-1}(R)(K)\text{ is dense in }(\pi\circ\rho^{-1}\circ p)^{-1}(R)\}$ is dense in $\mathbb{P}^{1}(K)$, thus $\mathcal{C}(K)$ is dense in $\mathcal{C}$. ∎
@@
The morphism $\phi:\mathcal{C}\rightarrow S$ has geometrically integral generic fiber.
@@
Assume that $\phi$ does not have a geometrically integral generic fiber; then the same holds for $\phi^{\text{n}}$, and we get a relative normalization decomposition:
@@
First extension. The map $p$ admits a rational section $S\dashrightarrow\mathcal{C},\ P\mapsto(P,Q)$ with $Q\in D_{P}$ such that Q=[2]PQ=[-2]P on the elliptic curve $D_{P}$ with origin $\mathcal{O}$ (see Remark 3.6). For general $P$, this point $Q$ has multiplicity $1$ in the intersection $D_{P}\cap C_{P}$ and is thus smooth in $C_{P}$, hence $(P,Q)$ lies in the smooth locus of $\mathcal{C}$. In particular, this section lifts to a map $s:S\dashrightarrow\mathcal{C}^{\text{n}}$. The composition $\phi^{\text{n}}\circ s:S\dashrightarrow S$ is the map $2]:S\dashrightarrow S$, and thus $K(S^{\prime})$ is contained in the degree-$4$ extension $K(S)_{2}/K(S)$ (with $K(S)_{2}\cong K(S)$ as abstract fields) given by multiplication by [2][-2].
Second extension. For general $P$, the restriction of the normalization morphism $\nu:\mathcal{C}^{\text{n}}\to\mathcal{C}$ above $C_{P}\cong\{P\}\times C_{P}\subset\mathcal{C}$ is the normalization morphism $\nu_{P}:C_{P}^{\text{n}}\to C_{P}$ of $C_{P}$. Since $C_{P}$ has a simple node at $P$, the inverse image $\nu_{P}^{-1}(P)$ comprises two (reduced) geometric points. It follows that the inverse image $V$ in $\mathcal{C}^{\text{n}}$ of a sufficiently small non-empty Zariski-open $U$ of the diagonal $\Delta:=\{(P,P):P\in S^{o}\}\subset\mathcal{C}$ is an étale double cover $V\to U$. The composition
$V\to\mathcal{C}^{\text{n}}\xrightarrow{\phi^{\text{n}}}S$
coincides with the composition $V\to U\subset\Delta\cong S^{o}\subset S$. It is thus a generically finite map of degree $2$, and by (1) the field extension $K(S^{\prime})/K(S)$ of degree $\geq 2$ embeds in the étale extension $K(V)/K(S)$ of degree $2$. Hence $K(S^{\prime})=K(V)$.
Linear disjointness. Since the Mordell–Weil group of $\mathcal{E}$ is torsion-free [27, Thm. 7.4], the degree-$4$ extension $K(S)_{2}/K(S)$ given by multiplication by [2][-2] has no non-trivial proper subextensions, as any of these would have degree $2$ and would give rise to a cyclic 2-isogeny on the elliptic surface $\mathcal{E}$, hence to a non-trivial $2$-torsion section. This immediately implies linear disjointness and concludes the proof. ∎
@@
Consider the family of curves $\{C_{P}\}_{P\in S^{o}}$ in $S$, and take the corresponding family of inverse images $\{\rho^{-1}(C_{P})\}_{P\in S^{o}}$ in $\mathcal{E}$.
The theorem follows by applying Theorem 1.1 to $\mathcal{E}$ endowed with the fibration $\pi$ and the family $\{\rho^{-1}(C_{P})\}_{P\in S^{o}}$. We verify the hypotheses below.
Geometrically integral generic fiber. The total space for the family $\{\rho^{-1}(C_{P})\}_{P\in S^{o}}$ is the fiber product $\mathcal{C}^{\prime}:=\mathcal{C}\times_{\phi,S,\rho}\mathcal{E}$. The generic fiber of $\mathcal{C}^{\prime}\to\mathcal{E}$ is the same as the generic fiber of $\mathcal{C}\to S$, and this is geometrically integral by Proposition 3.12.
Simple connectedness. All the fibers of $\pi$ are geometrically integral (they are Weierstrass curves), hence the only irreducible divisors of $\mathcal{E}$ that are vertical for $\pi$ are precisely these fibers, i.e. the inverse images $\rho^{-1}(D)$ of the curves $D$ in the pencil $|{-K_{S}}|$. Proposition 3.7 implies that the general curve $C_{P}$ intersects all $D\in|{-K_{S}}|,D\neq\{w=0\}$ transversally , hence the general inverse image $\rho^{-1}(C_{P})$ intersects $\rho^{-1}(D)$ transversally for all such $D$. It follows that $F$ is empty, so $S\setminus F=S$ is simply connected.
Zariski-density. $\mathcal{C}(K)$ is dense in $\mathcal{C}$ by Proposition 3.11 (essentially by Nijgh’s work). Since $\mathcal{C}^{\prime}$ is birational to $\mathcal{C}$, we see that $\mathcal{C}^{\prime}(K)$ is dense in $\mathcal{C}^{\prime}$. ∎
@@
Let $U\subset\mathbb{A}_{K}^{9}$ be the non-empty Zariski-open whose $\bar{K}$-points are $9$-tuples $(a,b,\ldots,f_{3})$ corresponding to a smooth $S$, and let $X\subset U\times\mathbb{P}(2,3,1,1)$ be the total space associated to the parametric family of Definition 3.1. Then the projection $\pi=pr_{1}:X\to U$ is smooth surjective and its fibers are del Pezzo surfaces of degree $1$. Note that $X$ is birational to the affine hypersurface $Y$ in $\mathbb{A}_{K}^{11}$ defined by the equation
$y^{2}=x^{3}+(af(z)+b)x+(cf(z)^{2}+df(z)+e),$
with $f(t):=f_{3}t^{3}+f_{2}t^{2}+f_{1}t+f_{0}$. The hypersurface $Y$ is rational via projection to any coordinate appearing linearly, e.g. $e$.
@@
For any $x\in V_{N}(K)$, the del Pezzo surface $\pi^{-1}(\pi(x))$ has the Hilbert property by Theorem 1.2. Now, the morphism $\pi|_{V_{N}}:V_{N}\to U$ has geometrically integral generic fiber and thus sends non-thin sets to non-thin sets. In particular, since $V_{N}(K)$ is non-thin by Hilbert’s Irreducibility Theorem (recall that $X$, and thus $V_{N}$, is rational), the set $\pi(V_{N}(K))$ is non-thin.